# Uncertainty Modeling in Graph Neural Networks via Stochastic Differential Equations

### v1

@@ 2 Methodology
Figure 1: The diagram illustrates the evolution of input data in the latent space through an SDE, showing the discretization of a sample path (purple line) at three timesteps with corresponding vector fields and graph embeddings. Initially, nodes representing two classes (white and orange) are poorly separated. However, as the model progresses, aggregating features and propagating information through the graph, the vector field helps make the class boundaries more distinct.
@@ 2.1 Model Definition
LGNSDEs are designed to capture the stochastic latent evolution of $\mathbf{H}(t)$ on graph-structured data. We use an Ornstein-Uhlenbeck (OU) prior process, represented by:
@@
where we set the drift and diffusion functions, $\mathbf{F}_{\mathcal{G}}$ and $\mathbf{G}_{\mathcal{G}}$, to constants and consider them hyperparameters. Moreover, $d\mathbf{W}(t)$ is a Wiener process. The approximate posterior is defined as:
@@
Let $\mathbf{Y}$ be a collection of target variables, e.g., class labels, for some of the graph nodes. Given $\mathbf{Y}$ we train our model with variational inference, with the ELBO computed as:
@@
where $\mathbf{H}(t_{0})$ are the node-wise features $\mathbf{X}_{\text{in}}$ in the graph $\mathcal{G}$. In practice, this is not feasible since the posterior drift $\mathbf{F}_{\mathcal{G},\phi}$ is parametrised by a neural network. We numerically solve this integral with a standard Stochastic Runge-Kutta method (Rößler, 2010). We then use a Monte Carlo approximation to get the expectation of $\mathbf{H}(t_{0})$ and approximate the posterior predictive distribution as:
@@ 3 Theoretical Guarantees
In this section, we present key results on the stability and robustness of our framework under mild assumptions (Appendix A.2). Firstly, we address the fundamental question of whether our proposed models provide meaningful uncertainties. By showing that the variance of the model output is bounded by the variance of the latent representation, we highlight the ability of LGNSDEs to capture and quantify the inherent uncertainty in the system. We formalize our findings and summarize them in the following lemma:
@@
Under assumptions 1-3, there exists a unique strong solution to a LGNSDE of the form:
@@
Under assumptions 1-3, consider two initial conditions $\mathbf{H}_{0}$ and $\tilde{\mathbf{H}}_{0}=\mathbf{H}_{0}+\delta\mathbf{H}(0)$, where $\delta\mathbf{H}(0)\in\mathbb{R}^{n\times d}$ is a small perturbation in the initial node features with $\|\delta\mathbf{H}(0)\|_{F}=\epsilon$. Assume that $\mathbf{H}_{0}$ is taken from a compact set $\mathcal{H}\subseteq\mathbb{R}^{n\times d}$. Then, the deviation between the solutions $\mathbf{H}(t)$ and $\tilde{\mathbf{H}}(t)$ of the LGNSDE with these initial conditions remains bounded across time $t$, specifically:
@@ 4 Experiments
We evaluate LGNSDEs on 5 datasets (see B and C for details on these datasets and hyperparameters), we compare it to GNODE (Poli et al., 2019), GCN Kipf and Welling (2016), Bayesian GCN (BGCN) (Hasanzadeh et al., 2020), and an ensemble of GCNs (Lin et al., 2022).
@@ A.3 Proofs
TODO: Mention the subtleties of having latent matrices instead of latent vectors.
@@ Appendix C Additional Results
LGNSDE (ours) 0.9667 ± 0.0036 0.9111 ± 0.0072 0.9691 ± 0.0032 0.9909 ± 0.0004 0.9007 ± 0.0091
AURC (↓) GCN 0.9963 ± 0.0010 0.9959 ± 0.0014 0.9994 ± 0.0003 0.9990 ± 0.0010 0.9994 ± 0.0005
GNODE nan ± nan 0.9967 ± 0.0008 0.9996 ± 0.0002 0.9996 ± 0.0002 0.9933 ± 0.0130
BGCN 0.9968 ± 0.0008 0.9963 ± 0.0005 0.9993 ± 0.0002 0.9988 ± 0.0005 0.9997 ± 0.0000
Ensemble 0.9971 ± 0.0007 0.9967 ± 0.0010 0.9993 ± 0.0002 0.9990 ± 0.0004 0.9987 ± 0.0026
LGNSDE (ours) 0.9968 ± 0.0003 0.9967 ± 0.0006 0.9994 ± 0.0003 0.9991 ± 0.0006 0.9995 ± 0.0003

### v3

@@ 2 Methodology
Figure 1: The diagram shows the evolution of input data in latent space $\mathbf{H}(t)$ through an SDE, with sample paths (purple) and confidence bands representing variance. At three timesteps, we visualize graph embeddings, where nodes (white and orange) become more separable over time due to the influence of the vector field. The inset axes represent latent dimensions, while the purple and yellow background highlights the magnitude and direction of the vector field guiding the latent dynamics.
@@ 2.1 Model Definition
LGNSDEs are designed to capture the stochastic latent evolution of $\mathbf{H}(t)$ on graph-structured data. We use an Ornstein-Uhlenbeck (OU) prior process, represented by
@@
where we set the drift and diffusion functions, $\mathbf{F}_{\mathcal{G}}$ and $\mathbf{G}_{\mathcal{G}}$, to constants and consider them hyperparameters. Moreover, $d\mathbf{W}(t)$ is a Wiener process. The approximate posterior is defined as
@@
Let $\mathbf{Y}$ be a collection of target variables, e.g., class labels, for some of the graph nodes. Given $\mathbf{Y}$ we train our model with variational inference, with the ELBO computed as
@@
where $\mathbf{H}(t_{0})$ are the node-wise features $\mathbf{X}_{\text{in}}$ in the graph $\mathcal{G}$. In practice, this is not feasible since the posterior drift $\mathbf{F}_{\mathcal{G},\phi}$ is parametrised by a neural network. We numerically solve this integral with a standard Stochastic Runge-Kutta method (Rößler, 2010). We then use a Monte Carlo approximation to get the expectation of $\mathbf{H}(t_{0})$ and approximate the posterior predictive distribution as
@@ 3 Theoretical Guarantees
In this section, we present key results on the stability and robustness of our framework under mild assumptions (Appendix A.2). Firstly, we address the fundamental question of whether our proposed models provide meaningful uncertainties. By showing that the variance of the latent representation bounds the model output variance, we highlight the ability of LGNSDEs to capture and quantify inherent uncertainty in the system. The latent representation is the underlying structure from which the model’s output is generated, i.e. the uncertainty in the latent space directly influences the uncertainty in predictions. We formalize this in the following lemma:
@@
Under assumptions 1-3, there exists a unique strong solution to an LGNSDE of the form
@@
Under assumptions 1-3, consider two initial conditions $\mathbf{H}_{0}$ and $\tilde{\mathbf{H}}_{0}=\mathbf{H}_{0}+\delta\mathbf{H}(0)$, where $\delta\mathbf{H}(0)\in\mathbb{R}^{n\times d}$ is a small perturbation in the initial node features with $\|\delta\mathbf{H}(0)\|_{F}=\epsilon$. Assume that $\mathbf{H}_{0}$ is taken from a compact set $\mathcal{H}\subseteq\mathbb{R}^{n\times d}$. Then, the deviation between the solutions $\mathbf{H}(t)$ and $\tilde{\mathbf{H}}(t)$ of the LGNSDE with these initial conditions remains bounded across time $t$, specifically
@@ 4 Experiments
We evaluate LGNSDEs on 5 datasets (see A.2 for details on these datasets and hyperparameters), we compare it to GNODE (Poli et al., 2019), GCN Kipf and Welling (2016), Bayesian GCN (BGCN) (Hasanzadeh et al., 2020), and an ensemble of GCNs (Lin et al., 2022).
@@ Appendix B Details of The Experimental Setup
LGNSDE (Our) 0.9667 ± 0.0036 0.9111 ± 0.0072 0.9691 ± 0.0032 0.9909 ± 0.0004 0.9007 ± 0.0091
AURC (↓) GCN 0.9966 ± 0.0007 0.9966 ± 0.0011 0.9994 ± 0.0005 0.9987 ± 0.0015 0.9994 ± 0.0004
GNODE nan ± nan 0.9967 ± 0.0011 0.9994 ± 0.0004 0.9998 ± 0.0001 0.9915 ± 0.0163
BGCN 0.9972 ± 0.0004 0.9963 ± 0.0010 0.9994 ± 0.0002 0.9989 ± 0.0005 0.9996 ± 0.0004
ENSEMBLE 0.9970 ± 0.0012 0.9967 ± 0.0012 0.9994 ± 0.0002 0.9989 ± 0.0006 0.9996 ± 0.0005
LGNSDE (Our) 0.9970 ± 0.0003 0.9971 ± 0.0005 0.9995 ± 0.0003 0.9997 ± 0.0002 0.9995 ± 0.0005